### v1

# An Algebraic Roadmap of Particle Theories Part I: General construction

Abstract
It is a curious fact that our most well-studied BSM theories can inform us about the Standard Model itself. Just over a decade ago, [1], Baez and Huerta explained how the Standard Model’s gauge group can be seen to coincide with the intersection of Georgi and Glashow’s SU(5), and Pati and Salam’s $SU(4)\times SU(2)\times SU(2)$, inside Spin(10). More recently, [2], [3], it was shown how (up to a factor of B-L) the Spin(10) model, the Pati-Salam model, the Left-Right Symmetric model, the Standard Model pre-Higgs mechanism, and the Standard Model post-Higgs mechanism are interlinked via a sequence of division algebraic reflections. Between these three papers, [1], [2], [3], are six theories that have been studied extensively by particle physicists. In this article, we demonstrate how the full set of six familiar particle models may be laid out in one connected algebraic particle roadmap.
The inclusion of a quaternionic reflection within the network further differentiates $W^{\pm}$ bosons from the $Z^{0}$ boson in comparison to the Standard Model. It may introduce subtle new considerations for the phenomenology of electroweak symmetry breaking.
pacs:
112.10.Dm, 2.60.Rc, 12.38.-t, 02.10.Hh, 12.90.+b
I Introduction
You shall know a word by the company it keeps.
- J.R. Firth, 1957.
If Quantum Mechanics, [4] and Special Relativity, [5], have taught us anything, it is that objects are often best understood not in isolation, but rather, in relation to their peers. Perhaps the Standard Model of particle physics is no exception.
In this article, we identify a set of repeating connections, in series and in parallel, between a number of neighbouring particle models. Among these nine models, six are well-known. We pinpoint the post-Higgs Standard Model at the cluster’s most highly constrained corner.
Figure 1: Simplified preview of an algebraic particle roadmap. The detailed version appears in Figure (6) at the end of this article.
For a glance ahead, see Figure (1). There, the network we are about to assemble will be built up using the natural operator spaces of the division algebras $\mathbb{R},\mathbb{C},\mathbb{H},$ and $\mathbb{O}.$ These operator spaces (multiplication algebras) are readily described by Clifford algebras.
Since at least as early as the 1970s, physicists and mathematicians have been identifying Clifford algebraic and division algebraic patterns within the architecture of elementary particles. Now half a century ago, Günaydin and Gürsey proposed a quark model based on octonions, [6]. Shortly thereafter came a series of papers by Casalbuoni et al who proposed particle models based on a large variety of Clifford algebras, [7], [8], [9]. Since these days, many authors have invested a great deal of time and effort to the endeavour. In what is apologetically far from an exhaustive list, we point out the Clifford algebraic work of Trayling and Baylis (Cl(0,7)) [10], Barrett (via NCG) [11], Zenczykowski ($\mathbb{C}l(6)$) [12], Connes et al (via NCG) [13], Gording and Schmidt-May ($\mathbb{C}l(6)$) [14], Todorov ($\mathbb{C}l(10)$) [15], Borštnik (Cl(p,q)) [16], the division algebraic work Adler ($\mathbb{H}$) [17], De Leo ($\mathbb{H}$) [18], Toppan et al ($\mathbb{H},\mathbb{O}$) [19], Baez ($\mathbb{R},\mathbb{C},\mathbb{H}$) [20], Duff et al ($\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{O}$) [21], Hughes ($\mathbb{O}$) [22], Catto et al ($\mathbb{O}$) [23], Gresnigt ($\mathbb{O}$) [24], Asselmeyer-Maluga ($\mathbb{H},\mathbb{O}$) [25], Bolokhov ($\mathbb{H}$) [26], Vaibhav and Singh ($-\mathbb{H}$ and split-$\mathbb{O}$) [27], Boyle ($J_{3}(\mathbb{O})$) [28], Jackson ($\mathbb{O}$) [29], Hunt ($\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{O}$) [30], Lasenby ($\mathbb{O}$) [31], Manogue et al ($E_{8}$) [32], [33], Hun Jang, (Hypercomplex) [35], Hiley ($-\mathbb{H}$) [36], and most recently Penrose (split-$\mathbb{O}$) [34], the $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$ work of Dixon, [37], Perelman, [38], Chester et al, [39], Köplinger, [40].
This paper is the first of a series; see also [41], [42]. Although the model described here was non-trivial to find, readers may appreciate that a large number of its results can be confirmed easily.
II In context
The particle roadmap introduced in this article extends directly from previous findings of Baez and Huerta [1], and Furey and Hughes [2], [3]. We summarize these earlier findings here. This section provides the historical background for this article, but is not strictly necessary to understand its results.
II.1 Algebra of grand unified theories
It is a surprisingly little-known fact amongst particle physicists that many of our most well-studied theories are interrelated.
The birth of the Spin(10) model in the mid-1970s came only a few hours before that of Georgi and Glashow’s SU(5) model [43], [44]. On its own, Georgi and Glashow’s model posits a seemingly ad hoc fermionic particle content as the $\mathbf{10}\oplus\mathbf{5^{*}}$ of SU(5). Why, one might wonder, this curious combination of irreducible representations?
However, embedding these irreps inside the proposed $\mathbf{16}$ of Spin(10) justifies the representation structure, while adding in a sterile neutrino in the form of an $SU(5)$ singlet.
$\begin{array}[]{ccc}\mathfrak{so}(10)&\supset&\mathfrak{su}(5)\\ \mathbf{16}&\hskip 28.45274pt\mapsto&\mathbf{10}\oplus\mathbf{5^{*}}\oplus\mathbf{1}\end{array}$ (1)
Hence the Spin(10) model offers guidance for its younger SU(5) sibling.
In the same era, another (partially) unified theory was constructed by Pati and Salam, [45]. Their intention was to capitalize on an observed pattern-matching between quarks and leptons. Via the group $SU(4)\times SU(2)\times SU(2)$, the Standard Model’s three “red, green, blue” quarks colours were augmented to include a fourth “lilac” lepton colour. Furthermore, a symmetry between left- and right-handed particles was conjectured. Curiously enough, Pati and Salam’s model also fits neatly into the $\mathbf{16}$ of Spin(10):
$\begin{array}[]{ccc}\mathfrak{so}(10)&\supset&\mathfrak{su}(4)\oplus\mathfrak{su}(2)\oplus\mathfrak{su}(2)\\ \mathbf{16}&\hskip 28.45274pt\mapsto&\left(\mathbf{4},\mathbf{2},\mathbf{1}\right)\oplus\left(\mathbf{4^{*}},\mathbf{1},\mathbf{2}\right)\end{array}$ (2)
Again, we find that the Spin(10) model justifies for the Pati-Salam model an otherwise arbitrary choice in fermion representations.
Now, more surprising than the kinship between the Spin(10) and SU(5) models, or the kinship between Spin(10) and Pati-Salam models, is an unexpected kinship between SU(5), Pati-Salam, and the Standard Model. That is, in [1], Baez and Huerta report that the Standard Model’s gauge group coincides exactly with the intersection between SU(5) and Pati-Salam symmetries. The following Figure (2) relates these four well-studied particle models.
Figure 2: Baez and Huerta, [1], explain how Standard Model symmetries (SM) result from the intersection of Georgi and Glashow’s SU(5) and the Pati-Salam (PS) symmetries SU(4)$\times$×SU(2)$\times$×SU(2). Spin(10) may be seen to break to PS by preserving the splitting $\mathbf{10}\mapsto\mathbf{6}\oplus\mathbf{4}$ of SO(10)’s defining representation. Alternatively, Spin(10) may be seen to break to SU(5) by preserving a grading and volume form of the spinor space $\Lambda\mathbb{C}^{5}$.
But what could prompt the symmetries to break in this way? Reference [1] lists three independent conditions:
1. A requirement that the defining $\mathbf{10}$ of $\mathfrak{so}$(10) splits into $\mathbf{6}\oplus\mathbf{4}$, (this breaks $\mathfrak{so}(10)\mapsto\mathfrak{su}(4)\oplus\mathfrak{su}(2)\oplus\mathfrak{su}(2)$),
2. The preservation of a $\mathbb{Z}$-grading on the spinorial $\Lambda\mathbb{C}^{5}\simeq\mathbf{16}\oplus\mathbf{16^{*}}$ of $\mathfrak{so}$(10), (this breaks $\mathfrak{so}(10)\mapsto\mathfrak{u}(5)$),
3. The invariance of a volume form in the spinorial $\Lambda\mathbb{C}^{5}\simeq\mathbf{16}\oplus\mathbf{16^{*}}$ of $\mathfrak{so}$(10), (this breaks $\mathfrak{u}(5)\mapsto\mathfrak{su}(5)$).
Here, $\Lambda\mathbb{C}^{5}$ denotes the exterior algebra generated by vectors of 5 complex dimensions. See also footnotes in [9].
From where do these conditions arise? Baez and Huerta encourage readers to decipher the meaning behind these mysterious constraints.
II.2 Division algebraic symmetry breaking
That the first constraint,
$\begin{array}[]{ccc}\mathfrak{so}(10)&\mapsto&\mathfrak{so}(6)\oplus\mathfrak{so}(4)\\ \mathbf{10}&\hskip 28.45274pt\mapsto&\mathbf{6}\oplus\mathbf{4}\end{array}$ (3)
derives from the octonions was first discovered by Furey and Hughes.
These results were distributed widely amongst colleagues on the 16th of February, 2021, [2], [3], and discussed extensively with some of them, [46]. In particular, Kirill Krasnov contributed to our efforts by following suit and reproducing our results, thus confirming them.
This octonionic structure may be seen to be responsible for sending the Spin(10) model to the Pati-Salam model, and the SU(5) model to the Standard Model.
As will be explained in detail later on in this text, the Spin(10) $\mapsto$ Pati-Salam and SU(5) $\mapsto$ Standard Model transitions occur upon the requirement that the symmetries be invariant under a certain type of octonionic reflection. Please see Figure (3). Going beyond, it was found in [2], [3] that invariance under a complementary octonionic reflection furthermore sends this Pati-Salam model ($\mathfrak{so}(6)\oplus\mathfrak{so}(4)=\mathfrak{su}(4)\oplus\mathfrak{su}(2)\oplus\mathfrak{su}(2)$) to the Left-Right Symmetric model ($\mathfrak{su}(3)\oplus\mathfrak{su}(2)\oplus\mathfrak{su}(2)\oplus\mathfrak{u}(1)$). Then, invariance under a quaternionic reflection sends the Left-Right Symmetric model to the Standard Model, augmented with a B-L symmetry ($\mathfrak{su}(3)\oplus\mathfrak{su}(2)\oplus\mathfrak{u}(1)\oplus\mathfrak{u}(1)$). Finally, [2], invariance under a complex conjugate reflection sends the pre-Higgs Standard Model ($\mathfrak{su}(3)\oplus\mathfrak{su}(2)\oplus\mathfrak{u}(1)$) to the post-Higgs Standard Model ($\mathfrak{su}(3)\oplus\mathfrak{u}(1)$).
In short, invariance under $\mathbb{O},$ $\mathbb{H},$ and $\mathbb{C}$ reflections links five well-established particle models in a cascade of breaking symmetries.
Figure 3: References [2], [3] introduced a cascade of breaking symmetries. The Spin(10) model breaks to the Pati-Salam model, (PS), via the octonions. The Pati-Salam model breaks to the Left-Right Symmetric model, (LR), again via the octonions. The Left-Right Symmetric model breaks to the Standard model + B-L, (SM+), via the quaternions. The Standard model + B-L breaks to the Standard Model’s unbroken symmetries, together with B-L, (UB+), via the complex numbers. Two open questions remained in this model: (1) Why the lingering B-L symmetry? (2) Could there be a more natural way to introduce chirality in the quaternionic “$\mathbb{H}_{\textup{R}}$” step described in [3]?
However, as with [1], some open questions remain. Namely, in the quaternionic Left-Right Symmetric $\mapsto$ Standard Model step,
1. An additional B-L symmetry persisted,
2. Chirality was introduced by hand.
Could it be possible to tame these unruly features?
In this article, we demonstrate that both of these issues may be addressed simultaneously. Furthermore, we offer a concise answer to those challenges posed by Baez and Huerta as to where from their mysterious conditions may be seen to originate.
As will be seen, the overarching theme may be described as the implementation of simultaneous group actions on the fermionic representations. These simultaneous group actions, or more precisely, Lie algebra actions, allow us to merge the diagrams of Figures 2 and 3. The result (Figure 6) is a detailed roadmap of particle models.
In order to keep the discussion as concise as possible, we will work on the level of Lie algebras. Those interested in group theoretic details are encouraged to consult [1].
II.3 A consistency condition
It is hard not to notice that, when augmented by a sterile neutrino, the fermionic representations of the Standard Model materialize as broken pieces of the $\mathbf{16}$ spinor representation of Spin(10). However, the $\mathfrak{su}(3)\oplus\mathfrak{su}(2)\oplus\mathfrak{u}(1)$ Standard Model symmetries acting on this representation space are unmistakably unitary. Could there be more than one type of symmetry constraint working in concert?
On a not-obviously related front, let us point out the minimal left ideal construction of spinors. As will be explained in detail shortly, such spinors may be constructed from the $2n$ generating $\Gamma$-matrices of certain $\mathbb{C}l(2n)$ Clifford algebras. Collectively, linear combinations of these $\Gamma$-matrices form the defining $2n$ vector representation of the Clifford algebra’s $\mathfrak{so}$(2n) symmetries. But this then begs the question: If the spinor is constructed directly from these vectors, should it not also transform under a group action induced by them?
In this article, we entertain the idea that Standard Model fermions should simultaneously transform under both the known $\mathfrak{spin}(2n)$ spinor action, and also under the known $\mathfrak{so}(2n)$ multivector action. This consistency condition reduces $\mathfrak{spin}(2n)$ to $\mathfrak{su}(n)$. In one stroke, this single requirement addresses both the chirality and B-L issues plaguing [3]. It also fuses Baez and Huerta’s conditions 2. and 3. in Section II.1 into a single constraint. It may be seen to be algebraically equivalent to a statement made in a footnote of [9], where a volume form is fixed. Here, we derive the result instead as a consistency condition in the construction of fermions as multivectors.
The main results of these papers were first made public in an abstract circulated for the Perimeter Institute Octonions and Standard Model conference in 2021. We include the abstract here: [47].
II.4 Triality’s example
As an aside, we mention that the idea of simultaneous group actions is not new to representation theory, [22], [33]. A phenomenon known as octonionic triality has three copies of $\mathbb{O}$ each transforming under Spin(8). Distinct $\mathfrak{so}$(8) actions on these 8D spaces allow one copy to be identified with the $\psi$ spinor, another with the $\widetilde{\psi}$ conjugate spinor, and the final copy with the vector representation, $V$. Now, introducing a requirement that each of the corresponding group actions coincide then reduces the original $\mathfrak{so}$(8) symmetry to $\mathfrak{g}_{2}$. While $\mathfrak{so}$(8) generated the triality symmetry of the octonions, $\mathfrak{g}_{2}$ generates their automorphisms.
Parallels between automorphisms embedded inside triality symmetries, and the standard model internal symmetries embedded inside $\mathfrak{spin}(10)$ will be made in the third paper of this series, [42].
III Simultaneous group actions of the multivector type
In this section, we offer an explanation for the breaking of symmetry from $\mathfrak{spin}(2n)\mapsto\mathfrak{su}(n)$, first proposed in [47]. In known models, this corresponds to transitions
$\begin{array}[]{rcl}\textup{Spin(10)}&\mapsto&\textup{SU(5),}\\ \textup{Pati-Salam}&\mapsto&\textup{Standard Model,}\\ \textup{Left-Right Symmetric}&\mapsto&\textup{Standard Model.}\end{array}$
While algebraic single-step methods breaking $\mathfrak{spin}(2n)\mapsto\mathfrak{u}(n)$ are ubiquitous in the literature, algebraic single-step methods breaking $\mathfrak{spin}(2n)\mapsto\mathfrak{su}(n)$ are not.
III.1 Three types of spinor
We argue in this subsection that not all spinor constructions are created equal. We compare three inequivalent constructs with increasing complexity.
(I) The most common description of a spinor in the literature is simply as a column vector (equivalently, single-indexed tensor) with entries in $\mathbb{R}$ or $\mathbb{C}$.
(II) As an alternative to column vectors, certain spinors may be constructed by upgrading these vector spaces to exterior algebras, $\Lambda\mathbb{C}^{n}$. In this case, the spinor is endowed with a $\mathbb{Z}$-grading and a concept of a wedge product on its elements.
(Aside) It has long been known that this construction leads to a complementary description in terms of differential forms and vector fields, eg [9]. Namely,
$a_{j}=\textup{d}x^{j}\wedge\hskip 28.45274pta_{j}^{\dagger}=i\frac{\partial}{\partial x^{j}},$ (4)
where the $\{a_{j}\}$ represent the $n$ generating elements of $\Lambda\mathbb{C}^{n}$.
Augmenting spinors-as-column-vectors to spinors-as-exterior-algebras introduces a structural richness that would otherwise be invisible in the more common column vector approach. However, this is not the end of the line.
(III) The construction of spinors as minimal left ideals (MLI) takes this $\Lambda\mathbb{C}^{n}$, and adds another layer of structure beyond it. As will be described shortly, this extra structure appears in the form of a non-trivial vacuum state. The MLI spinor construction leads to a more unified description by embedding the spinor directly into the Clifford algebra. We now introduce this construction to readers not already familiar.
III.2 Minimal left ideal construction
Consider a Clifford algebra, $\mathbb{C}l(2n)$, over the complex numbers, and generated by basis vectors $\Gamma_{k}$ for $k\in\{1,\dots,2n\}$, and $n\in\mathbb{N}>0$. Let $\dagger$ denote an anti-linear involution such that $(\Gamma_{k_{1}}\Gamma_{k_{2}})^{\dagger}=\Gamma_{k_{2}}^{\dagger}\Gamma_{k_{1}}^{\dagger}$. In this setup, we are free to choose $\Gamma_{k}$ such that $\Gamma_{k}^{\dagger}=-\Gamma_{k},$ and
$\{\Gamma_{k_{1}},\Gamma_{k_{2}}\}:=\Gamma_{k_{1}}\Gamma_{k_{2}}+\Gamma_{k_{2}}\Gamma_{k_{1}}=-2\delta_{k_{1}k_{2}},$ (5)
and we will indeed do so.
From here, we construct an alternative generating basis for $\mathbb{C}l(2n)$ as consisting of $n$ raising operators, $a_{j}$, and $n$ lowering operators, $a^{\dagger}_{j}$, for $j\in\{1,2,\dots n\}$. (Naming conventions were chosen so that this work matches previous articles, [2], [3].) Explicitly,
$a_{j}:=\frac{1}{2}\left(-\Gamma_{j}+i\Gamma_{n+j}\right)\hskip 28.45274pta^{\dagger}_{j}:=\frac{1}{2}\left(\Gamma_{j}+i\Gamma_{n+j}\right).$ (6)
From these definitions and equations (5), we have that
$\begin{array}[]{l}\{a_{j_{1}},a_{j_{2}}\}=\{a_{j_{1}}^{\dagger},a_{j_{2}}^{\dagger}\}=0\\ \{a_{j_{1}},a^{\dagger}_{j_{2}}\}=\delta_{j_{1}j_{2}}\end{array}$ (7)
for $j_{1},j_{2}\in\{1,2,\dots n\}$. We define $\Omega:=a_{1}a_{2}\cdots a_{n}$. Then we form a hermitian vacuum state as
$v:=\Omega^{\dagger}\Omega=a_{n}^{\dagger}\cdots a_{2}^{\dagger}a_{1}^{\dagger}a_{1}a_{2}\cdots a_{n}.$ (8)
We are now ready to define a minimal left ideal, $\Psi,$ as
$\Psi:=\mathbb{C}l(2n)\hskip 1.42262ptv.$ (9)
These raising and lowering operators allow us to make apparent a $\mathbb{Z}$-grading within $\Psi$. It can be shown that
$\begin{array}[]{lc}\Psi=&c_{12\dots n}\hskip 1.42262pta_{1}a_{2}\cdots a_{n}v\\ &\vdots\\ &+\hskip 2.84526ptc_{12}\hskip 1.42262pta_{1}a_{2}v+c_{13}\hskip 1.42262pta_{1}a_{3}v\dots+c_{n-1\hskip 1.42262ptn}\hskip 1.42262pta_{n-1}a_{n}v\\ &+\hskip 2.84526ptc_{1}\hskip 1.42262pta_{1}v+c_{2}\hskip 1.42262pta_{2}v\dots+c_{n}\hskip 1.42262pta_{n}v\\ &+\hskip 2.84526ptc_{0}\hskip 1.42262ptv\end{array}$ (10)
for $c_{0},c_{j},\dots,c_{12\dots n}\in\mathbb{C}$. From here, it is straightforward to to see that $\Psi$ takes the form of a Fock space, familiar to physicists.
It can be shown that the omission of $v$ in equation (10) leads to an exterior algebra. Hence, the MLI construction may be seen to bestow upon an exterior algebra the extra structure of a non-trivial vacuum state.
III.3 Adjoint, vector, and spinor representations
It is known, [48], that the Lie algebra $\mathfrak{so}(2n)$ may be represented by bivectors $r_{k_{1}k_{2}}\Gamma_{k_{1}}\Gamma_{k_{2}}$ for $r_{k_{1}k_{2}}\in\mathbb{R}$. Here, $k_{1},k_{2}\in\{1,2,\dots 2n\}$, and $k_{1}\neq k_{2}$. Multiplication is given by the commutator,
$[r_{k_{1}k_{2}}\Gamma_{k_{1}}\Gamma_{k_{2}},\hskip 2.84526ptr_{k_{3}k_{4}}\Gamma_{k_{3}}\Gamma_{k_{4}}].$
Similarly, the defining $2n$-dimensional vector representation of $\mathfrak{so}(2n)$ may be written as $V:=\sum_{k=1}^{2n}V_{k}\hskip 1.42262pt\Gamma_{k}$ for $V_{k}\in\mathbb{R}$. Its infinitesimal transformations under $\mathfrak{so}(2n)$ are then given by
$\delta V=[r_{k_{1}k_{2}}\Gamma_{k_{1}}\Gamma_{k_{2}},\hskip 1.42262ptV].$ (11)
It is important to note that in this formalism, basis vectors are viewed as carrying the transformations, not the coefficients.
[…]
$\delta\Psi=r_{k_{1}k_{2}}\Gamma_{k_{1}}\Gamma_{k_{2}}\Psi.$ (12)
III.4 $\mathfrak{spin}(2n)$ $\mapsto$ $\mathfrak{su}(n)$ via the
multivector condition
In equation (10), we explained that spinors can be constructed purely from the generating $\{a_{1},a_{2},\dots a_{1}^{\dagger},a_{2}^{\dagger},\dots\}$ of the Clifford algebra. Using equation (6), we may then express this spinor purely in terms of the $\{\Gamma_{k}\}$. Explicitly, these minimal left ideals are composed of multivectors
$\Psi=c_{0}+c_{i_{1}}\Gamma_{i_{1}}+c_{i_{2}i_{3}}\Gamma_{i_{2}}\Gamma_{i_{3}}+\dots,$ (13)
for $i_{1},i_{2},i_{3},\dots\in\{1,2,\dots 2n\}$ and for some $c_{0},c_{i_{1}},c_{i_{2}i_{3}},\dots$ $\in\mathbb{C}$.
As mentioned in equation (11), these $\{\Gamma_{k}\}$ form the vector representation $V:=\sum_{k=1}^{2n}V_{k}\hskip 1.42262pt\Gamma_{k}$ of $\mathfrak{so}(2n)$,
$\delta V=[r_{k_{1}k_{2}}\Gamma_{k_{1}}\Gamma_{k_{2}},\hskip 1.42262ptV].$ (14)
From this vantage point, it is then natural to wonder: Shouldn’t the generating vectors induce a transformation rule on the minimal left ideals that were built from them?
Equation (14) supplies a derivation. One would expect an induced infinitesimal transformation of $\Psi$ to go as
[…]
which simplifies to
$\delta\Psi=[r_{k_{1}k_{2}}\Gamma_{k_{1}}\Gamma_{k_{2}},\hskip 1.42262pt\Psi]\hskip 2.84526pt.$ (16)
However, we already defined the infinitesimal transformation of spinors in equation (12) as
$\delta\Psi=r_{k_{1}k_{2}}\Gamma_{k_{1}}\Gamma_{k_{2}}\Psi.$ (17)
So, which shall it be: equation (16) or equation (17)?
Let us entertain the idea that both symmetry actions (16) and (17) be simultaneously obeyed. We will refer to this consistency condition as the multivector condition.
Readers may confirm that only an $\mathfrak{su}(n)$ subalgebra of $\mathfrak{so}(2n)$ survives the multivector condition.
$\begin{array}[]{c}\delta\Psi=[r_{k_{1}k_{2}}\Gamma_{k_{1}}\Gamma_{k_{2}},\hskip 1.42262pt\Psi]\hskip 2.84526pt=r_{k_{1}k_{2}}\Gamma_{k_{1}}\Gamma_{k_{2}}\Psi\\ \Rightarrow\\ \mathfrak{so}(2n)\hskip 5.69054pt\mapsto\hskip 5.69054pt\mathfrak{su}(n).\end{array}$ (18)
Explicitly, a generic element of the surviving $\mathfrak{su}(n)$ subalgebra may be written in terms of raising and lowering operators as
$\begin{array}[]{lll}\ell_{n}&=&R_{j_{1}j_{2}}\left(a_{j_{1}}a^{\dagger}_{j_{2}}-a_{j_{2}}a^{\dagger}_{j_{1}}\right)\\ &+&R_{j_{1}j_{2}}^{\prime}i\left(a_{j_{1}}a^{\dagger}_{j_{2}}+a_{j_{2}}a^{\dagger}_{j_{1}}\right)\\ &+&R_{j}\hskip 1.42262pti\left(a_{j}a^{\dagger}_{j}-a_{j+1}a^{\dagger}_{j+1}\right),\end{array}$ (19)
where $j_{1}\neq j_{2}$ and $R_{j_{1}j_{2}},R_{j_{1}j_{2}}^{\prime},R_{j}\in\mathbb{R}$. In terms of $\{\Gamma_{j}\}$, these same $\mathfrak{su}(n)$ elements may be represented as
$\begin{array}[]{lll}\ell_{n}&=&r_{j}\hskip 1.42262pt\Gamma_{j}\Gamma_{j+n}\\ &+&r_{j_{1}j_{2}}\left(\Gamma_{j_{1}}\Gamma_{j_{2}}+\Gamma_{j_{1}+n}\Gamma_{j_{2}+n}\right)\\ &+&r_{j_{1}j_{2}}^{\prime}\left(\Gamma_{j_{1}}\Gamma_{j_{2}+n}+\Gamma_{j_{2}}\Gamma_{j_{1}+n}\right),\end{array}$ (20)
where $r_{j},r_{j_{1}j_{2}},r_{j_{1}j_{2}}^{\prime}\in\mathbb{R},$ and $\sum_{j=1}^{n}r_{j}=0.$
III.5 Significance
This result is significant for a number of reasons.
(+) It shows that a single consistency condition, (18), can offer an answer to both of the open challenges 2. and 3. posed by Baez and Huerta in Section II.1.
(+) As will appear later in this text, it eliminates the persistent unwanted B-L symmetry from [3].
(+) We will also see that in the physically interesting cases of
$\begin{array}[]{rcl}\textup{Spin(10)}&\mapsto&\textup{SU(5),}\\ \textup{Pati-Salam}&\mapsto&\textup{Standard Model,}\\ \textup{Left-Right Symmetric}&\mapsto&\textup{Standard Model,}\end{array}$
this consistency condition supplies a natural explanation for the Standard Model’s infamous chirality.
IV Simultaneous group actions of the reflective type
IV.1 Motivation for considering $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$
Recent work, [2], [3], introduced the result that one generation of unconstrained fermions can be identified with one copy of $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$,
one $\textup{one generation}\hskip 5.69054pt\leftrightarrow\hskip 5.69054pt\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}.$
Furthermore, the division algebraic substructure of this algebra (surprisingly) led to a cascade of symmetry breakings in well-known particle models:
$\begin{array}[]{l}\textup{Spin(10)}\mapsto\textup{Pati-Salam}\mapsto\textup{Left-Right Symmetric}\mapsto\\ \hskip 21.33955pt\textup{Pre-Higgs Standard Model (+ B-L)}\mapsto\\ \hskip 42.67912pt\textup{Post-Higgs Standard Model (+ B-L).}\end{array}$
However, instead of identifying one generation with $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$, we will now construct one generation as a minimal left ideal within $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$’s multiplication algebra,
[…]
Doing so will allow us to implement the multivector condition from the previous section.
$\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$’s multiplication algebra, to be defined shortly, is isomorphic to its complex endomorphisms, $End\left(\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}\right)$. It is also isomorphic as a matrix algebra to the complex Clifford algebra $\mathbb{C}l(10)$. Hence it provides the facilities to build up a Spin(10) model, and due to its division algebraic substructure, the facilities to break that model down.
IV.2 The algebra $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$
The algebra $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$ is also known as the Dixon algebra, due to its independent implementation in early particle models by Dixon. Readers are encouraged to see [37] for an alternative perspective.
Throughout this article, all tensor products will be assumed to be over $\mathbb{R}$ unless otherwise stated. We write the standard $\mathbb{R}$-basis for $\mathbb{H}$ as $\{\epsilon_{0},\epsilon_{1},\epsilon_{2},\epsilon_{3}\}$, where $\epsilon_{0}=1$, $\epsilon_{j}^{2}=-1$ for $j\in\{1,2,3\}$, and $\epsilon_{1}\epsilon_{2}=\epsilon_{3}$, with cyclic permutations. These relations may be written more succinctly as $\epsilon_{i}\epsilon_{j}=-\delta_{ij}+\varepsilon_{ijk}\hskip 1.42262pt\epsilon_{k}$ for $i,j,k\in\{1,2,3\}$, where $\varepsilon_{ijk}$ is the usual totally anti-symmetric tensor with $\varepsilon_{123}=1$.
Similarly, we write the standard $\mathbb{R}$-basis for $\mathbb{O}$ as $\{e_{0},e_{1},\dots,e_{7}\}$, where $e_{0}=1$ and $e_{i}e_{j}=-\delta_{ij}+f_{ijk}\hskip 1.42262pte_{k}$ for $i,j,k\in\{1,2,\dots 7\}$. Here, $f_{ijk}$ is a totally anti-symmetric tensor with $f_{ijk}=1$ when $ijk\in\{124,235,346,457,561,672,713\}$. The remaining values of $f_{ijk}$ are determined by anti-symmetry, and vanish otherwise. We let $i$ denote a complex imaginary unit as usual.
Let $\mathbb{A}:=\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}=\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$. Note that $\mathbb{A}$ is naturally an algebra over $\mathbb{C}$: scalar multiplication by $c\in\mathbb{C}$ is defined by setting $c(x\otimes y\otimes w)=cx\otimes y\otimes w$ for $x\in\mathbb{C},y\in\mathbb{H},w\in\mathbb{O}$. Multiplication of elements is defined by setting $(x_{1}\otimes y_{1}\otimes w_{1})(x_{2}\otimes y_{2}\otimes w_{2})=x_{1}x_{2}\otimes y_{1}y_{2}\otimes w_{1}w_{2}$ for all $x_{1},x_{2}\in\mathbb{C},y_{1},y_{2}\in\mathbb{H}$, and $z_{1},z_{2}\in\mathbb{O}$. A $\mathbb{C}$-basis for $\mathbb{A}$ is given by $\{1\otimes\epsilon_{\mu}\otimes e_{\nu}\mid\mu\in\{0,1,2,3\},\nu\in\{0,1,\dots,7\}\}$. We see that $\mathbb{A}$ is a 32 $\mathbb{C}$-dimensional non-commutative, non-associative algebra.
From now on, we identify $\mathbb{C}$, $\mathbb{H}$, and $\mathbb{O}$ with their images in $\mathbb{A}$ under the natural inclusion maps, thus writing $c$ instead of $c\otimes\epsilon_{0}\otimes e_{0}$ for any $c\in\mathbb{C}$, writing $\epsilon_{\mu}$ for $1\otimes\epsilon_{\mu}\otimes e_{0}$, and writing $e_{\nu}$ for $1\otimes\epsilon_{0}\otimes e_{\nu}$ (here $\mu\in\{0,1,2,3\},\hskip 2.84526pt\nu\in\{0,1,\dots,7\}$). Arbitrary elements of $\mathbb{A}$ are then written as $\sum_{\mu,\nu}c_{\mu\nu}\epsilon_{\mu}e_{\nu}$, where $c_{\mu\nu}\in\mathbb{C},$ and we see that $\epsilon_{\mu}e_{\nu}=e_{\nu}\epsilon_{\mu}$ for all $\mu,\nu$.
For those less comfortable with formal definitions, an example may help. Suppose $a,b\in\mathbb{A}$ with $a=4\epsilon_{1}e_{2}$ and $b=\left(5+i\right)\epsilon_{2}e_{4}$. Then $ab=\left(20+4i\right)\epsilon_{3}e_{1}.$
IV.3 From $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$ to its space of linear operators
In works dating back as early as [6], it was clear that the $\mathfrak{su}(3)$ adjoint part of gluons would be destined to reside not in $\mathbb{O},$ but rather in the space of linear operators on $\mathbb{O}$. This space of linear operators can be realized as $\mathbb{A}$’s multiplication algebra.
If bosons are to reside within $\mathbb{A}$’s multiplication algebra, then why not include fermions there as well? One advantage of including fermions within $\mathbb{A}$’s multiplication algebra is that, unlike with $\mathbb{A}$, this space is now large enough to comfortably accommodate three generations. Hence, we will position these fermions within $\mathbb{A}$’s multiplication algebra, in the form of minimal left ideals.
First let us define what we mean by multiplication algebra. Suppose $x,y$ are elements in an algebra $\mathbb{D}.$ We then define $L_{x}(y):=xy$. Similarly, define $R_{x}(y):=yx$. Both $L_{x}$ and $R_{x}$ may then be seen as (possibly distinct) linear maps sending $y$ to some new element in $\mathbb{D}.$ Hence, $L_{x}$ and $R_{x}$ $\in End(\mathbb{D}).$
For $x,y,z\in\mathbb{D},$ these linear maps may be composed as in $L_{x}\circ L_{y}(z)=x(y(z))$, $R_{x}\circ R_{y}(z)=((z)y)x,$ $L_{x}\circ R_{y}(z)=x((z)y),$ $R_{x}\circ L_{y}(z)=(y(z))x$, etc. Composition of these maps, $\circ$, defines an associative multiplication rule. Furthermore, maps may be added in the obvious way. For example, $(L_{x}+L_{y})(z)=xz+yz.$
We define the left multiplication algebra of $\mathbb{D}$ as the subalgebra of $End(\mathbb{D})$ generated by the $\{L_{y}\hskip 1.42262pt|\hskip 1.42262pty\in\mathbb{D}\}$. Similarly, the right multiplication algebra of $\mathbb{D}$ is the subalgebra of $End(\mathbb{D})$ generated by $\{R_{y}\hskip 1.42262pt|\hskip 1.42262pty\in\mathbb{D}\}$. We denote the left multiplication algebra of an algebra $\mathbb{D}$ as $\mathcal{L}_{\mathbb{D}},$ and its right multiplication algebra as $\mathcal{R}_{\mathbb{D}}.$
As an example, consider the multiplication of two elements $A,B\in End(\mathbb{A})$ with $A=4L_{\epsilon_{1}e_{2}}$ and $B=\left(5+i\right)L_{\epsilon_{2}e_{4}}$. Then $AB:=A\circ B=\left(20+4i\right)L_{\epsilon_{1}e_{2}}\circ L_{\epsilon_{2}e_{4}}=\left(20+4i\right)L_{\epsilon_{3}e_{2}}\circ L_{e_{4}}\neq\left(20+4i\right)L_{\epsilon_{3}e_{1}}.$
IV.4 Clifford factors
In this subsection, we examine the substructure of $\mathbb{A}$’s multiplication algebra. We will find that it is this substructure that ultimately leads the Spin(10) model to splinter into some of its broken successors.
What are the multiplication algebras of $,\mathbb{C},$$\mathbb{H},$ and $\mathbb{O}$? Of $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$? It can be confirmed that the following isomorphisms hold (as matrix algebras):
$\begin{array}[]{lllll}&&\mathcal{L}_{\mathbb{C}}=\mathcal{R}_{\mathbb{C}}&\hskip 5.69054pt\leftrightarrow&Cl\left(0,1\right)\\ &&\mathcal{L}_{\mathbb{H}}&\hskip 5.69054pt\leftrightarrow&Cl\left(0,2\right)\\ &&\mathcal{R}_{\mathbb{H}}&\hskip 5.69054pt\leftrightarrow&Cl\left(0,2\right)\\ &&\mathcal{L}_{\mathbb{O}}\simeq\mathcal{R}_{\mathbb{O}}&\hskip 5.69054pt\leftrightarrow&Cl\left(0,6\right).\\ \end{array}$ (21)
Since $\mathbb{C}$ is abelian, $\mathcal{L}_{\mathbb{C}}$ and $\mathcal{R}_{\mathbb{C}}$ are equal elementwise. That is, $L_{c}=R_{c}\hskip 5.69054pt\forall c\in\mathbb{C}$. We identify these linear maps with $Cl(0,1),$ where the vector generating $Cl(0,1)$ is given by multiplication by the complex imaginary unit, $L_{i}=R_{i}$.
Unlike the complex numbers, $\mathbb{H}$ is non-abelian, and perhaps unsurprisingly, $\mathcal{L}_{\mathbb{H}}$ and $\mathcal{R}_{\mathbb{H}}$ typically provide distinct linear maps on $\mathbb{H}$. We may identify each of $\mathcal{L}_{\mathbb{H}}$ and $\mathcal{R}_{\mathbb{H}}$ with one copy of $Cl(0,2)$, thereby identifying the combined multiplication algebras of $\mathbb{H}$ with $Cl\left(0,2\right)\otimes Cl\left(0,2\right).$ For concreteness, we will take $L_{\epsilon_{1}}$ and $L_{\epsilon_{2}}$ to generate the first copy of $Cl(0,2),$ and $R_{\epsilon_{1}}$ and $R_{\epsilon_{2}}$ to generate the second copy of $Cl(0,2),$ although there is clearly a continuum of equivalent choices.
Finally, the left- and right-multiplication algebras of $\mathbb{O}$ may each be identified with the Clifford algebra $Cl(0,6)$. However, unlike with the quaternions, it is possible to show that each element of $\mathcal{R}_{\mathbb{O}}$ gives the same linear map on $\mathbb{O}$ as some element in $\mathcal{L}_{\mathbb{O}}.$ For example, $\forall f\in\mathbb{O}$,
$\begin{array}[]{lll}R_{e_{7}}f:=fe_{7}&=&\frac{1}{2}\left(-e_{7}f+e_{1}\left(e_{3}f\right)+e_{2}\left(e_{6}f\right)+e_{4}\left(e_{5}f\right)\right)\\ &=&\frac{1}{2}\left(-L_{e_{7}}+L_{e_{1}}L_{e_{3}}+L_{e_{2}}L_{e_{6}}+L_{e_{4}}L_{e_{5}}\right)f,\end{array}$ (22)
where $e_{1},e_{2},\dots e_{7}$ represent octonionic imaginary units. Therefore we see that although $L_{a}$ and $R_{a}$ are not equal for every $a\in\mathbb{O}$, $\mathcal{L}_{\mathbb{O}}$ and $\mathcal{R}_{\mathbb{O}}$ do provide the same set of linear maps on $\mathbb{O}.$ Hence, we write that $\mathcal{L}_{\mathbb{O}}\simeq\mathcal{R}_{\mathbb{O}}.$ The set of octonionic linear maps, identified with $Cl(0,6),$ may be generated by $\{L_{e_{j}}\}$, or equivalently by $\{R_{e_{j}}\},$ where $j=1,\dots,6.$ Again, we emphasize that a continuum of equivalent choices exists (6-sphere $S^{6}$) .
In short, we find that the linear maps coming from the multiplication algebras of $,\mathbb{C},$$\mathbb{H},$ and $\mathbb{O}$ are given by $Cl(0,1),$ $Cl(0,2)\otimes Cl(0,2),$ and $Cl(0,6)$ respectively. The gradings for these Clifford algebras can result from choosing the Clifford algebra’s generators to be multiplication by an orthogonal set of imaginary units. $Cl(0,1)$ may be generated by $L_{i}=R_{i};$ the two factors of $Cl(0,2)\otimes Cl(0,2)$ may be generated by $\{L_{\epsilon_{1}},L_{\epsilon_{2}}\}$ and $\{R_{\epsilon_{1}},R_{\epsilon_{2}}\};$ $Cl(0,6)$ may be generated by $\{L_{e_{j}}\}$, or alternatively by $\{R_{e_{j}}\},$ where $j\in\{1,\dots,6\}.$
Finally, combining $\mathcal{L}_{\mathbb{C}}=\mathcal{R}_{\mathbb{C}},$ $\mathcal{L}_{\mathbb{H}},$ and $\mathcal{L}_{\mathbb{O}}\simeq\mathcal{R}_{\mathbb{O}},$ gives $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$’s left multiplication algebra as
$\begin{array}[]{l}\mathcal{L}_{\mathbb{C}}\otimes\mathcal{L}_{\mathbb{H}}\otimes\mathcal{L}_{\mathbb{O}}\\ \simeq Cl(0,1)\otimes Cl(0,2)\otimes Cl(0,6)\simeq\mathbb{C}l(8).\end{array}$ (23)
Combining $\mathcal{L}_{\mathbb{C}}=\mathcal{R}_{\mathbb{C}},$ $\mathcal{L}_{\mathbb{H}},$ $\mathcal{R}_{\mathbb{H}},$ and $\mathcal{L}_{\mathbb{O}}\simeq\mathcal{R}_{\mathbb{O}},$ gives $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$’s full multiplication algebra as
$\begin{array}[]{l}\mathcal{L}_{\mathbb{C}}\otimes\mathcal{L}_{\mathbb{H}}\otimes\mathcal{R}_{\mathbb{H}}\otimes\mathcal{L}_{\mathbb{O}}\\ \simeq Cl(0,1)\otimes Cl(0,2)\otimes Cl(0,2)\otimes Cl(0,6)\simeq\mathbb{C}l(10).\end{array}$ (24)
IV.5 Hermitian conjugation
We define hermitian conjugation, $\dagger$, on elements of $End(\mathbb{A})$ as the involution such that $L_{e_{j}}^{\dagger}=-L_{e_{j}}$ $\forall j\in\{1,2,\dots 7\}$, $L_{\epsilon_{m}}^{\dagger}=-L_{\epsilon_{m}}$ and $R_{\epsilon_{m}}^{\dagger}=-R_{\epsilon_{m}}$ $\forall m\in\{1,2,3\}$, and $L_{i}^{\dagger}=-L_{i}.$ As with matrix algebras, the hermitian conjugate defined here obeys $\left(ab\right)^{\dagger}=b^{\dagger}a^{\dagger}$ $\forall a,b\in End\left(\mathbb{A}\right).$
IV.6 Division algebraic reflections
We have now explored the substructure of $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}$’s multiplication algebra. Next, we would like to know how one goes about putting this substructure to use.
In short, we will employ the $\mathbb{Z}_{2}$-gradings that are inherent to each Clifford algebraic factor in equation (24). Or, said another way, we will make use of generalized notions of reflection.
Referring to the first factor in equation (24), we know that $\mathcal{L}_{\mathbb{C}}$ is isomorphic to the Clifford algebra $Cl(0,1)$. In this case, complex conjugation, denoted $*_{\overrightarrow{\mathbb{C}}}\vspace{.5mm}$, maps $L_{i}\mapsto-L_{i},$ and defines the $\mathbb{Z}_{2}$-grading on this Clifford algebra. It also induces a reflection of any complex number across the real axis.
It is straightforward to identify analog reflections in the subsequent Clifford algebraic factors of equation (24). For example, a $\mathbb{Z}_{2}$-grading on $\mathcal{L}_{\mathbb{H}}\simeq Cl(0,2)$ may be induced by a quaternionic reflection *$*_{\overrightarrow{\mathbb{H}}}:L_{\epsilon_{m}}\mapsto-L_{\epsilon_{m}}$ for $m\in\{1,2\}$. Similarly, a $\mathbb{Z}_{2}$-grading on $\mathcal{L}_{\mathbb{O}}\simeq Cl(0,6)$ may be induced by an octonionic reflection *$*_{\overrightarrow{\mathbb{O}}}:L_{e_{j}}\mapsto-L_{e_{j}}$ for $j\in\{1,2,\dots 6\}$.
Certain attentive readers might rightfully question why these octonionic and quaternionic reflections involve only six and two imaginary units respectively. Why not seven and three? This comes from the left multiplication algebras being isomorphic to $Cl(0,6)$ and $Cl(0,2)$, not $Cl(0,7)$ and $Cl(0,3)$. Nonetheless, rest assured that the final imaginary units become reintegrated when these reflections are composed with Clifford algebraic reversion: $(\Gamma_{i}\Gamma_{j})^{\rho}=\Gamma_{j}\Gamma_{i}.$ It is tedious, but straightforward, to confirm that $L_{e_{7}}=L_{e_{1}}L_{e_{2}}L_{e_{3}}L_{e_{4}}L_{e_{5}}L_{e_{6}}$. Hence, $(L_{e_{7}}^{*_{\overrightarrow{\mathbb{O}}}})^{{\rho}}=L_{e_{7}}^{{\rho}}=-L_{e_{7}}.$ Similarly, $L_{\epsilon_{3}}=L_{\epsilon_{1}}L_{\epsilon_{2}}$. Hence, $(L_{\epsilon_{3}}^{*_{\overrightarrow{\mathbb{H}}}})^{{\rho}}=L_{\epsilon_{3}}^{{\rho}}=-L_{\epsilon_{3}}.$
For ease of exposition, we have chosen specific imaginary units above. However, readers should be aware that an entire 6-sphere of equivalent choices exist for $\mathcal{L}_{\mathbb{O}}$, a 2-sphere of equivalent choices exist for $\mathcal{L}_{\mathbb{H}}$, and a 0-sphere ($\mathbb{Z}_{2}$) of equivalent choices exists for $\mathcal{L}_{\mathbb{C}}$.
IV.7 An explicit division algebraic representation of $\mathbb{C}l(10)$
There are many ways to generate $\mathbb{C}l(10)\simeq End(\mathbb{A})$ so that the results in this paper materialize. One such set is given by
$\Gamma_{j}=iL_{e_{j}}L_{e_{7}}R_{\epsilon_{1}},\hskip 14.22636pt\Gamma_{m+6}=L_{e_{7}}L_{\epsilon_{m}}R_{\epsilon_{1}},\hskip 14.22636pt\Gamma_{10}=R_{\epsilon_{2}}$ (25)
for $j\in\{1,\dots 6\},$ and $m\in\{1,2,3\}.$ (Another, possibly simpler, starting point has been found recently by John Barrett). We define
$\Gamma_{11}:=\prod_{p=1}^{10}\Gamma_{p},$ (26)
which in this case becomes
$\Gamma_{11}=-R_{\epsilon_{3}}.$ (27)
It allows us to define a chirality operator $-i\Gamma_{11}=iR_{\epsilon_{3}}$. We specify raising operators $a_{c}$ and lowering operators $a_{c}^{\dagger}$ as
$\begin{array}[]{lll}a_{1}^{\dagger}:=\frac{1}{2}(\Gamma_{5}+i\Gamma_{4})&&a_{1}:=\frac{1}{2}(-\Gamma_{5}+i\Gamma_{4})\\ a_{2}^{\dagger}:=\frac{1}{2}(\Gamma_{3}+i\Gamma_{1})&&a_{2}:=\frac{1}{2}(-\Gamma_{3}+i\Gamma_{1})\\ a_{3}^{\dagger}:=\frac{1}{2}(\Gamma_{6}+i\Gamma_{2})&&a_{3}:=\frac{1}{2}(-\Gamma_{6}+i\Gamma_{2})\\ a_{4}^{\dagger}:=\frac{1}{2}(\Gamma_{8}+i\Gamma_{7})&&a_{4}:=\frac{1}{2}(-\Gamma_{8}+i\Gamma_{7})\\ a_{5}^{\dagger}:=\frac{1}{2}(\Gamma_{10}+i\Gamma_{9})&&a_{5}:=\frac{1}{2}(-\Gamma_{10}+i\Gamma_{9}),\\ \end{array}$ (28)
where readers may notice that we have made a permissible relabeling of indices relative to equation (6) in order to connect with previous work, [2], [3]. Defining a nilpotent object
$\Omega:=\prod_{c=1}^{5}a_{c}$ (29)
then sets our hermitian vacuum state as $v:=\Omega^{\dagger}\Omega$. Finally, we construct a minimal left ideal as
$\Psi:=\mathbb{C}l(10)v.$ (30)
This gives a $32\hskip 1.42262pt\mathbb{C}$ dimensional subspace of $\mathbb{C}l(10)$. For this first article in the series, we will be particularly interested in the $16\hskip 1.42262pt\mathbb{C}$ semi-spinor given by
$\Psi_{\textup{L}}:=\frac{1}{2}(1+i\Gamma_{11})\Psi.$ (31)
In the second article of this series, [41], we will then extend the semi-spinor in two inequivalent ways.
IV.8 Cascade of particle symmetries:
the $\mathfrak{spin}$(10) line
With our fermions now embedded within $End(\mathbb{A})$ we will demonstrate how division algebraic reflections prompt them to fragment. Please see Figure (4).
Figure 4: Four types of division algebraic reflections are relevant in the left-multiplication space. Invariance under the first octonionic reflection restricts Spin(10) to Pati-Salam (PS). Invariance under the second octonionic reflection restricts Pati-Salam to Left-Right Symmetric (LR). Invariance under a quaternionic reflection restricts Left-Right Symmetric to the Standard Model’s unbroken symmetries, together with a $\mathfrak{u}(1)$ related to B-L, and another $\mathfrak{u}(1)$ related to the $Z^{0}$ boson, (UBZ+). Invariance under complex conjugation eliminates the $\mathfrak{u}(1)$ related to the $Z^{0}$, leaving the Standard Model’s unbroken symmetries, together with B-L, (UB+). Readers can confirm that certain paths bifurcate.
A generic element of the 45-dimensional $\mathfrak{spin}$(10) may be written as real linear combinations of bivectors:
$\begin{array}[]{ll}\ell_{10}^{D}:=&r_{ij}L_{e_{i}}L_{e_{j}}+r_{m}L_{\epsilon_{m}}+r^{\prime}_{m}L_{\epsilon_{m}}L_{e_{7}}R_{\epsilon_{3}}\\ &+\hskip 2.84526ptr_{mj}^{\prime}iL_{\epsilon_{m}}L_{e_{j}}+r_{j}^{\prime\prime}iL_{e_{j}}L_{e_{7}}R_{\epsilon_{3}}\end{array}$ (32)
for $i,j\in\{1,\dots 6\}$ with $i\neq j,$ $m\in\{1,2,3\},$ and $r_{ij},r_{m},r^{\prime}_{m},r_{mj}^{\prime},r_{j}^{\prime\prime}\in\mathbb{R}$. Here the ‘D’ in $\ell_{10}^{D}$ refers to ‘Dirac’.
When applied to our semi-spinor $\Psi_{\textup{L}}$, this $\mathfrak{spin}$(10) action simplifies considerably:
$\begin{array}[]{lc}\ell_{10}^{D}\Psi_{\textup{L}}=&(r_{ab}L_{e_{a}}L_{e_{b}}+r^{\prime}_{mn}L_{\epsilon_{m}}L_{\epsilon_{n}}+r_{ma}^{\prime\prime}iL_{\epsilon_{m}}L_{e_{a}})\Psi_{\textup{L}},\\ &\mathfrak{spin}(10)\end{array}$ (33)
where this time $a,b\in\{1,\dots 7\}$ with $a\neq b,$ and $m,n\in\{1,2,3\}$ with $m\neq n,$ while $r_{ab},r^{\prime}_{mn},r_{ma}^{\prime\prime}\in\mathbb{R}.$ This Weyl action matches that found earlier in [3]. Let us then define
$\begin{array}[]{lc}\ell_{10}:=&r_{ab}L_{e_{a}}L_{e_{b}}+r^{\prime}_{mn}L_{\epsilon_{m}}L_{\epsilon_{n}}+r_{ma}^{\prime\prime}iL_{\epsilon_{m}}L_{e_{a}}.\\ &\mathfrak{spin}(10)\end{array}$ (34)
With this representation of $\mathfrak{spin}(10)$ at our disposal, we may now make use of the division algebraic reflections introduced in Section IV.6. That these reflections would reduce the SO(10) grand unified theory to other well-known particle models came originally as a surprise.
It is important to note that the $\mathfrak{spin}(10)$ element $\ell_{10}$ lives exclusively in the left-multiplication algebra of $\mathbb{A}$. Therefore there are four types of division algebraic reflections that will be relevant for us: those related to $\langle 1\rangle$ $\mathcal{L}_{\mathbb{O}},$ $\langle 2\rangle$ $\mathcal{R}_{\mathbb{O}},$ $\langle 3\rangle$ $\mathcal{L}_{\mathbb{H}},$ and $\langle 4\rangle$ $\mathcal{L}_{\mathbb{C}}.$
$\langle 1\rangle$ First of all, we would like to know: Which Lie subalgebra of our $\mathfrak{spin}(10)$ symmetries is invariant under an octonionic reflection? Readers can easily confirm that setting
${\ell}_{10}\Psi_{\textup{L}}={\ell}_{10}^{*_{\overrightarrow{\mathbb{O}}}}\Psi_{\textup{L}}\hskip 28.45274pt\forall\Psi_{L}$ (35)
restricts $\mathfrak{spin}(10)$’s $\ell_{10}$ to the Pati-Salam symmetries, $\ell_{\textup{PS}}:$
$\begin{array}[]{rll}{\ell}_{\textup{PS}}\Psi_{\textup{L}}=(r_{ij}L_{e_{i}}L_{e_{j}}&+&r_{m}L_{\epsilon_{m}}+r^{\prime}_{m}iL_{\epsilon_{m}}L_{e_{7}})\Psi_{\textup{L}}\\ \mathfrak{spin}(6)&\oplus&\mathfrak{spin}(4)\end{array}$ (36)
for $i,j\in\{1,\dots 6\}$ with $i\neq j,$ and $m\in\{1,2,3\}.$ Note that $\mathfrak{spin}(6)=\mathfrak{su}(4),$ and $\mathfrak{spin}(4)=\mathfrak{su}(2)\oplus\mathfrak{su}(2)$.
Next, we will consider those symmetries furthermore invariant under an analogous octonionic reflection, deriving from right multiplication. Namely *$*_{\overleftarrow{\mathbb{O}}}:R_{e_{j}}\mapsto-R_{e_{j}}$ $\forall j\in\{1,\dots,6\}.$
$\langle 2\rangle$ Which Lie subalgebra of the Pati-Salam symmetries is invariant under this complementary octonionic reflection? With a little work, readers will find the Lie algebra of the Left-Right Symmetric model. Setting
${\ell}_{\textup{PS}}\Psi_{\textup{L}}={\ell}_{\textup{PS}}^{*_{\overleftarrow{\mathbb{O}}}}\Psi_{\textup{L}}\hskip 28.45274pt\forall\Psi_{L}$ (37)
restricts $\ell_{\textup{PS}}$ to the Left-Right Symmetric model’s gauge symmetries, $\ell_{\textup{LR}}:$
$\begin{array}[]{rll}{\ell}_{\textup{LR}}\Psi_{\textup{L}}=(r_{ij}^{\prime\prime}L_{e_{i}}L_{e_{j}}&+&r_{m}L_{\epsilon_{m}}+r^{\prime}_{m}iL_{\epsilon_{m}}L_{e_{7}})\Psi_{\textup{L}},\\ \mathfrak{u}(3)&\oplus&\mathfrak{spin}(4)\end{array}$ (38)
where the coefficients $r_{ij}^{\prime\prime}\in\mathbb{R}$ are this time restricted to give a $\mathfrak{u}(3)$ subalgebra of $\mathfrak{spin}(6)$. Explicitly, we have eight $\mathfrak{su}(3)_{C}$ generators
$\begin{array}[]{lll}i\Lambda_{1}:=\frac{1}{2}\left(L_{34}-L_{15}\right)&&i\Lambda_{2}:=\frac{1}{2}\left(L_{14}+L_{35}\right)\\ i\Lambda_{3}:=\frac{1}{2}\left(L_{13}-L_{45}\right)&&i\Lambda_{4}:=-\frac{1}{2}\left(L_{25}+L_{46}\right)\\ i\Lambda_{5}:=\frac{1}{2}\left(L_{24}-L_{56}\right)&&i\Lambda_{6}:=-\frac{1}{2}\left(L_{16}+L_{23}\right)\\ i\Lambda_{7}:=-\frac{1}{2}\left(L_{12}+L_{36}\right)&&i\Lambda_{8}:=\frac{-1}{2\sqrt{3}}\left(L_{13}+L_{45}-2L_{26}\right),\end{array}$ (39)
where $L_{ij}$ is shorthand for the octonionic $L_{e_{i}}L_{e_{j}}$. The generator for the $\mathfrak{u}(1)_{B-L}$ subalgebra of $\mathfrak{u}(3)$ may be described as
$i(B-L):=\frac{1}{3}(L_{13}+L_{26}+L_{45}).$ (40)
An alternative route was found by Boyle in [28] that maps Spin(10) directly to the Left-Right Symmetric model via *$*_{\overleftarrow{\mathbb{O}}}$ in the context of the exceptional Jordan algebra. This result likewise holds in our current model.
$\langle 3\rangle$ Next, fixing
${\ell}_{\textup{LR}}\Psi_{\textup{L}}={\ell}_{\textup{LR}}^{*_{\overrightarrow{\mathbb{H}}}}\Psi_{\textup{L}}\hskip 28.45274pt\forall\Psi_{L}$ (41)
restricts ${\ell}_{\textup{LR}}$ to a $\mathfrak{u}(3)\oplus\mathfrak{u}(1)\oplus\mathfrak{u}(1)$ subalgebra, which we will describe by elements ${\ell}_{{\textup{UBZ+}}}$:
$\begin{array}[]{rllrll}{\ell}_{{\textup{UBZ+}}}\Psi_{\textup{L}}=(r_{ij}^{\prime\prime}L_{e_{i}}L_{e_{j}}&+&r_{3}L_{\epsilon_{3}}&+&r^{\prime}_{3}iL_{\epsilon_{3}}L_{e_{7}})\Psi_{\textup{L}}.\\ \mathfrak{u}(3)&\oplus&\mathfrak{u}(1)&\oplus&\mathfrak{u}(1)\end{array}$ (42)
This less recognizable Lie subalgebra describes the Standard Model’s unbroken gauge symmetries, together with a B-L symmetry, and an additional $\mathfrak{u}(1)$ symmetry corresponding to the $Z^{0}$ boson.
$\langle 4\rangle$ Finally, requiring this action’s invariance under complex conjugation,
${\ell}_{{\textup{UBZ+}}}\Psi_{\textup{L}}={\ell}_{{\textup{UBZ+}}}^{*_{\overrightarrow{\mathbb{C}}}}\Psi_{\textup{L}}\hskip 28.45274pt\forall\Psi_{L}$ (43)
restricts ${\ell}_{{\textup{UBZ+}}}$ to a $\mathfrak{u}(3)\oplus\mathfrak{u}(1)$ Lie subalgebra, described by elements ${\ell}_{{\textup{UB+}}}$:
$\begin{array}[]{rll}{\ell}_{{\textup{UB+}}}\Psi_{\textup{L}}=(r_{ij}^{\prime\prime}L_{e_{i}}L_{e_{j}}&+&r_{3}L_{\epsilon_{3}})\Psi_{\textup{L}},\\ \mathfrak{u}(3)&\oplus&\mathfrak{u}(1)\end{array}$ (44)
thereby eliminating the symmetry corresponding to the $Z^{0}$ boson. This leaves us with the Standard Model’s unbroken gauge symmetries, after the Higgs mechanism, in addition to a B-L symmetry.
IV.9 Cascade of particle symmetries: the Georgi-Glashow line
We will now mark a path of broken symmetries, parallel to that of the previous section. This time, however, we will begin with Georgi and Glashow’s SU(5) model. Please see Figure (5).
Figure 5: The same division algebraic reflections as Figure (4) may be applied to the Georgi-Glashow model. We find that invariance under either of the octonionic reflections sends SU(5) to the Standard Model (SM). Unlike with [3], though, chirality now arises organically, and the unwanted B-L symmetry is eliminated. From here, invariance under a quaternionic reflection sends the Standard Model’s symmetries to the Standard Model’s unbroken symmetries, together with a $\mathfrak{u}(1)$ related to the $Z^{0}$ boson, (UBZ). Invariance under the complex reflection finally eliminates this $\mathfrak{u}(1)$ factor to give the Standard Model’s familiar unbroken symmetries, (UB). Again, complex conjugation provides a bypass that maps SM directly to UB.
The 24  dimensional $\mathfrak{su}$(5) Lie algebra may be represented as
$\begin{array}[]{lll}\ell_{5}&=&R_{cd}\left(a_{c}a^{\dagger}_{d}-a_{d}a^{\dagger}_{c}\right)\\ &+&R_{cd}^{\prime}\hskip 1.42262pti\left(a_{c}a^{\dagger}_{d}+a_{d}a^{\dagger}_{c}\right)\\ &+&R_{c}\hskip 1.42262pti\left(a_{c}a^{\dagger}_{c}-a_{c+1}a^{\dagger}_{c+1}\right),\end{array}$ (45)
for $c,d\in\{1,\dots 5\}$ and $c\neq d$. Here, $R_{cd},R_{cd}^{\prime},R_{c}\in\mathbb{R}$. This representation of $\mathfrak{su}(5)$ corresponds to the Lie subalgebra of $\mathfrak{spin}(10)$ surviving the multivector condition, as in equation (19).
Substituting in the division algebraic raising and lowering operators of equation (28) into equation (45) gives an $\mathfrak{su}(5)$ element, $\ell_{5}$, acting on $\Psi_{\textup{L}}$ as
$\begin{array}[]{lllll}&&\ell_{5}\Psi_{\textup{L}}=&&\\ &&\big{(}r_{j}i\Lambda_{j}+r^{\prime}_{m}L_{\epsilon_{m}}s+irY&&\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{1}\left(L_{\epsilon_{1}}L_{e_{4}}+L_{\epsilon_{2}}L_{e_{5}}\right)&+&it_{1}^{\prime}\left(L_{\epsilon_{1}}L_{e_{5}}-L_{\epsilon_{2}}L_{e_{4}}\right)\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{1}^{\prime\prime}\left(L_{\epsilon_{3}}L_{e_{4}}-iL_{e_{5}}L_{e_{7}}\right)&+&it_{1}^{\prime\prime\prime}(L_{\epsilon_{3}}L_{e_{5}}+iL_{e_{4}}L_{e_{7}})\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{2}\left(L_{\epsilon_{1}}L_{e_{1}}+L_{\epsilon_{2}}L_{e_{3}}\right)&+&it_{2}^{\prime}(L_{\epsilon_{1}}L_{e_{3}}-L_{\epsilon_{2}}L_{e_{1}})\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{2}^{\prime\prime}\left(L_{\epsilon_{3}}L_{e_{1}}-iL_{e_{3}}L_{e_{7}}\right)&+&it_{2}^{\prime\prime\prime}(L_{\epsilon_{3}}L_{e_{3}}+iL_{e_{1}}L_{e_{7}})\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{3}\left(L_{\epsilon_{1}}L_{e_{2}}+L_{\epsilon_{2}}L_{e_{6}}\right)&+&it_{3}^{\prime}(L_{\epsilon_{1}}L_{e_{6}}-L_{\epsilon_{2}}L_{e_{2}})\\ &&\hskip 1.42262pt+\hskip 1.99168ptit_{3}^{\prime\prime}\left(L_{\epsilon_{3}}L_{e_{2}}-iL_{e_{6}}L_{e_{7}}\right)&+&it_{3}^{\prime\prime\prime}(L_{\epsilon_{3}}L_{e_{6}}+iL_{e_{2}}L_{e_{7}})\big{)}\Psi_{\textup{L}}\\ &&\hskip 48.36958pt\mathfrak{su}(5)&&\end{array}$ (46)
where $j\in\{1,\dots 8\},$ $m\in\{1,2,3\}$, and $r_{j},r^{\prime}_{m},r,t_{k},t_{k}^{\prime},\vspace{.5mm}$ $t_{k}^{\prime\prime},t_{k}^{\prime\prime\prime}\vspace{.5mm}\in\mathbb{R}$. The idempotent $s$ is defined as $s:=\frac{1}{2}(1+iL_{e_{7}}),$ and preserves $\mathfrak{su}(2)_{\textup{L}}$-active states. The weak hypercharge generator, $iY$, is defined via
$iY\Psi_{\textup{L}}:=\left(\frac{1}{6}\left(L_{13}+L_{26}+L_{45}\right)-\frac{1}{2}L_{\epsilon_{3}}s^{*}\right)\Psi_{\textup{L}},$ (47)
following the conventions of [49].
With our $\mathfrak{su}(5)$ action now established, division algebraic reflections may be deployed so as to break these symmetries sequentially.
$\langle 1\rangle$ Which Lie subalgebra of Georgi and Glashow’s $\mathfrak{su}(5)$ exhibits immunity to an octonionic reflection? Readers may easily confirm that fixing the constraint
${\ell}_{5}\Psi_{\textup{L}}={\ell}_{5}^{*_{\overrightarrow{\mathbb{O}}}}\Psi_{\textup{L}}\hskip 28.45274pt\forall\Psi_{L}$ (48)
restricts $\mathfrak{su}(5)$ to the Standard Model’s symmetries, $\mathfrak{su}(3)\oplus\mathfrak{su}(2)\oplus\mathfrak{u}(1)$.
$\langle 2\rangle$ Interestingly, we find that the right-multiplication condition
${\ell}_{5}\Psi_{\textup{L}}={\ell}_{5}^{*_{\overleftarrow{\mathbb{O}}}}\Psi_{\textup{L}}\hskip 28.45274pt\forall\Psi_{L}$ (49)
has an identical effect. In either case, we are left with the Standard Model’s gauge symmetries acting on $\Psi_{\textup{L}}$ as
$\begin{array}[]{rccll}\ell_{\textup{SM}}\Psi_{\textup{L}}=\big{(}r_{j}i\Lambda_{j}&+&r^{\prime}_{m}L_{\epsilon_{m}}s&+&irY\big{)}\Psi_{\textup{L}}.\\ \mathfrak{su}(3)&\oplus&\mathfrak{su}(2)&\oplus&\mathfrak{u}(1)\end{array}$ (50)
It is worth taking a moment to point out that the result established here is not only obtaining the correct Standard Model Lie algebras. Over the years, a recurring challenge in algebraic models has been to also secure the correct chiral fermion representations. That the correct representations are indeed realized here will be established in [41].
$\langle 3\rangle$ It is perhaps this next quaternionic step that is of most phenomenological interest. The condition
${\ell}_{\textup{SM}}\Psi_{\textup{L}}={\ell}_{\textup{SM}}^{*_{\overrightarrow{\mathbb{H}}}}\Psi_{\textup{L}}\hskip 28.45274pt\forall\Psi_{L}$ (51)
can be seen to restrict $\mathfrak{su}(3)\oplus\mathfrak{su}(2)\oplus\mathfrak{u}(1)$ to $\mathfrak{su}(3)\oplus\mathfrak{u}(1)\oplus\mathfrak{u}(1)$. This resulting $\mathfrak{su}(3)\oplus\mathfrak{u}(1)\oplus\mathfrak{u}(1)$ action describes the Standard Model’s unbroken gauge symmetries, together with an additional $\mathfrak{u}(1)$ symmetry associated with the $Z^{0}$ boson. Explicitly, its action on $\Psi_{\textup{L}}$ reduces to
$\begin{array}[]{rccll}\ell_{\textup{UBZ}}\Psi_{\textup{L}}=\big{(}r_{j}i\Lambda_{j}&+&r^{\prime}_{3}L_{\epsilon_{3}}s&+&irY\big{)}\Psi_{\textup{L}}.\\ \mathfrak{su}(3)&\oplus&\mathfrak{u}(1)&\oplus&\mathfrak{u}(1)\end{array}$ (52)
Its meaning for electroweak symmetry breaking will be discussed briefly later in this text.
$\langle 4\rangle$ Finally we arrive at invariance under complex conjugation. Applying the constraint that
${\ell}_{\textup{UBZ}}\Psi_{\textup{L}}={\ell}_{\textup{UBZ}}^{*_{\overrightarrow{\mathbb{C}}}}\Psi_{\textup{L}}\hskip 28.45274pt\forall\Psi_{L}$ (53)
leaves us with the Standard Model’s unbroken gauge symmetries, $\mathfrak{su}(3)\oplus\mathfrak{u}(1)$. A generic element of this Lie algebra acts on $\Psi_{\textup{L}}$ as
$\begin{array}[]{lrcl}\ell_{\textup{UB}}\Psi_{\textup{L}}=&\big{(}r_{j}i\Lambda_{j}&+&r^{\prime}iQ\big{)}\Psi_{\textup{L}},\\ &\mathfrak{su}(3)&\oplus&\mathfrak{u}(1)\end{array}$ (54)
where $r^{\prime}\in\mathbb{R},$ and $iQ$ turns out to be none other than the electric charge generator,
$iQ:=\frac{1}{6}\left(L_{13}+L_{26}+L_{45}\right)-\frac{1}{2}L_{\epsilon_{3}}.$ (55)
Readers should take note that it is also possible to bypass the $*_{\overrightarrow{\mathbb{H}}}$ step in $\langle 3\rangle$. In this case, one may move from the Standard Model’s pre-Higgs $\mathfrak{su}(3)\oplus\mathfrak{su}(2)\oplus\mathfrak{u}(1)$ directly to the Standard Model’s post-Higgs $\mathfrak{su}(3)\oplus\mathfrak{u}(1)$, simply via $*_{\overrightarrow{\mathbb{C}}}\vspace{1mm}$. Again, this outcome originally came as a surprise.
Particularly satisfying about this last result is the fact that the familiar complex conjugate is what dictates the Standard Model’s final unbroken gauge symmetries. Conceivably, the Standard Model’s gauge group could have broken in a different way, had it been paired with a different Higgs sector. Could the complex conjugate ultimately be stewarding the symmetry breaking process?
V Roadmap
Now that we have charted out certain paths connecting familiar particle models, we will consolidate them into one detailed roadmap.
In this paper, we have listed five different forms of symmetry breaking steps. Namely, the multivector condition, $\Psi_{V}$, invariance under two types of octonionic reflection, *${*_{\overrightarrow{\mathbb{O}}}}$ and *$*_{\overleftarrow{\mathbb{O}}}$, one quaternionic reflection, $*_{\overrightarrow{\mathbb{H}}},\vspace{.5mm}$ and one complex reflection, $*_{\overrightarrow{\mathbb{C}}}$. These four types of reflection are precisely those that can be realized in terms of the left-multiplication algebra of $\mathbb{R}\otimes\mathbb{C}\otimes\mathbb{H}\otimes\mathbb{O}.$
The quaternionic reflection $*_{\overleftarrow{\mathbb{H}}},\vspace{.5mm}$ related to right multiplication, is isolated in the sense that, unlike the others, it is not expressible in terms of the left multiplication algebra. In the second article of this series, [41], we will see how it plays an important role in the description of spacetime symmetries.
Our culminating diagram results when all five symmetry breaking steps are amalgamated ($\Psi_{V}$, *${*_{\overrightarrow{\mathbb{O}}}}$, *$*_{\overleftarrow{\mathbb{O}}}$, $*_{\overrightarrow{\mathbb{H}}},$$*_{\overrightarrow{\mathbb{C}}}$). For reflections, we begin with the octonions, continue to the quaternions, and close with the complex numbers. Please see Figure (6) at the end of this article.
Figure 6: Six well-known models of elementary particles are interconnected algebraically into a detailed particle roadmap. The “SO(10)” model (Spin(10)), the Georgi-Glashow model (SU(5)), the Pati-Salam model (PS), the Left-Right Symmetric model (LR), the Standard Model pre-Higgs-mechanism (SM), and the Standard Model post-Higgs-mechanism (UB) are each interrelated via certain algebraic constraints. The multivector condition is represented by edges labeled as $\Psi_{V}$. Division algebraic reflection constraints are represented by edges labeled as *$*_{\mathbb{D}}$. Beyond these six well-known models, there are three more. For the definition of UBZ+ see equation (42); for the definition of UB+ see equation (44); for the definition of UBZ see equation (52).
From the outset, readers may notice an interesting feature of this 5-step network. Namely, not all pathways are unique; a number of them bifurcate.
As a first example, we identify two compatible symmetry breaking pathways from the Spin(10) model to the Left-Right Symmetric model. One may first restrict Spin(10) to those symmetries invariant under the octonionic reflection *$*_{\overrightarrow{\mathbb{O}}}$, resulting in the Pati-Salam model. Subsequently imposing invariance under *${*_{\overleftarrow{\mathbb{O}}}}$ then results in the Left-Right Symmetric model. As an alternate route, readers can bypass directly to the Left-Right Symmetric model by simply restricting Spin(10) once via *$*_{\overleftarrow{\mathbb{O}}}$, as in [28].
Of more immediate experimental relevance is the bifurcated symmetry breaking step from the Standard Model symmetries pre-Higgs mechanism, $\mathfrak{su}(3)_{\textup{C}}\oplus\mathfrak{su}(2)_{\textup{L}}\oplus\mathfrak{u}(1)_{\textup{Y}}$, to the Standard Model symmetries post-Higgs mechanism, $\mathfrak{su}(3)_{\textup{C}}\oplus\mathfrak{u}(1)_{\textup{Q}}$. Readers will encounter a direct path via $*_{\overrightarrow{\mathbb{C}}}\vspace{.5mm}$ from $\mathfrak{su}(3)_{\textup{C}}\oplus\mathfrak{su}(2)_{\textup{L}}\oplus\mathfrak{u}(1)_{\textup{Y}}\mapsto\mathfrak{su}(3)_{\textup{C}}\oplus\mathfrak{u}(1)_{\textup{Q}}$. This algebraic transition coincides with the Standard Model’s famous spontaneous symmetry breaking. On the other hand, these same endpoints may be realized by first imposing an intermediate constraint under $*_{\overrightarrow{\mathbb{H}}}$. This extra step acts to further break $W^{\pm}$ bosons away from the $Z^{0}$ relative to the Standard Model. It will be interesting to investigate whether or not such an intermediate symmetry breaking step could alter the $W^{\pm}$ mass prediction, relative to that of the Standard Model [54], [55].
VI Conclusion
In this article, we set out to understand the Standard Model of particle physics in relation to a number of neighbouring particle models. We summarize our findings:
In what may be viewed as a consistency condition, we introduced what we call the multivector condition. This constraint dictates that minimal left ideal fermions transform in accordance with the multivectors that comprise them. As a result, for internal symmetries, we find that the adjoint, vector, and spinor representations now each transform infinitesimally under the commutator. The minimal left ideal spinors that we constructed constitute a special case where the commutator and left multiplication coincide. The condition is algebraically equivalent to, but conceptually distinct from, constraints of earlier authors fixing the volume form of a $\Lambda\mathbb{C}^{5}$ spinor.
The multivector constraint, [47], put an end to the lingering B-L and chirality issues that once haunted [3]. It also offered one way to understand and consolidate the Spin(10) $\mapsto$ SU(5) symmetry breaking steps explained in [1].
The findings of Baez and Huerta, [1], and Furey and Hughes, [2], [3], are fused into a single particle roadmap. A total of nine particle models are shown to be interlinked via the multivector condition, and via constraints of invariance under generalized division algebraic reflections. Six of these models are familiar to particle physicists, namely the “SO(10)” model, the Georgi-Glashow model, the Pati-Salam model, the Left-Right Symmetric model, the Standard Model pre-Higgs mechanism, and the Standard Model post-Higgs mechanism.
The algebraic relations between Standard Model pre- and post-Higgs may be of special interest to phenomenologists. That is, we find it fortunate that the complex conjugate singles out precisely those $\mathfrak{su}(3)_{\textup{C}}\oplus\mathfrak{u}(1)_{\textup{Q}}$ gauge symmetries found at low energies, including correct assignments of electric charge, [41]. We find it interesting that an alternative symmetry breaking path to this same endpoint may exist, bypassed, and perhaps overshadowed, by the Higgs’ direct route. Could such a parallel path carry with it phenomenological implications for electroweak physics?
Acknowledgements.
These manuscripts have benefitted from numerous discussions with Beth Romano. The author is furthermore grateful for feedback and encouragement from John Baez, Sukṛti Bansal, John Barrett, Latham Boyle, Hilary Carteret, Mia Hughes, Kaushlendra Kumar, Agostino Patella, Shadi Tahvildar-Zadeh, Carlos Tamarit, Andreas Trautner, and Jorge Zanelli. This work was graciously supported by the VW Stiftung Freigeist Fellowship, and Humboldt-Universität zu Berlin.